OpenStudy (anonymous):

What is the slope of a line that passes through the points (2,8) and (6,12) A. -5 B. -1 C. 1 D. -5

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (2,8) x1 = 2 and y1 = 8 (6,12) x2 = 6 and y2 = 12 now sub slope = (12 - 8) / (6 - 2) slope = 4/4 which equals 1

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

texaschic101 Thank you for helping me, I just didn't really understand how to do slope and how to set it up! Now I can figure out how to do the next questions!!! Thanks!!!!!!!!

OpenStudy (texaschic101):

glad to help :)